OpenStudy (anonymous):

Calculus HW! please helps! Use properties of logs to simplify the expression

OpenStudy (anonymous):

\[\log9(x−\sqrt{x ^{2}-45)}+\log_{9} (x+\sqrt{x ^{2}-45})\]

OpenStudy (anonymous):

\[ \log a + \log b = \log \left( a\cdot b\right)\]

OpenStudy (anonymous):

Base is same so they can be multiplied just aspgpilot326 has written

OpenStudy (anonymous):

\[\left( a-b \right)\left( a+b \right)=a ^{2}-b ^{2}\]

OpenStudy (anonymous):

@surjithayer i did and I got

OpenStudy (anonymous):

i see nothing written